OpenStudy (anonymous):

Factor completely: 35x2 -x - 6

OpenStudy (anonymous):

(ax+b) * (cx+d) = 35x^2-x-6

OpenStudy (anonymous):

bd add up to -1 multiply to -6

OpenStudy (anonymous):

(ax+b) * (cx+d)=a*x^2*c+a*x*d+b*c*x+b*d=35x^2-x-6 (5*x+2)*(7*x-3)